### v2

# DecoHD: Decomposed Hyperdimensional Classification under Extreme Memory Budgets

Abstract
[…]
I Introduction
Hyperdimensional computing (HDC), rooted in vector-symbolic architectures (VSA), has gained traction as a lightweight and noise-tolerant paradigm for classification and learning on constrained platforms [14, 12, 9, 38, 37, 35, 36, 32]. By encoding data into dense high-dimensional hypervectors and manipulating them with simple algebraic operations, HDC classifiers combine attractive traits: inherent robustness to device-level non-idealities, parallelism amenable to hardware acceleration, and compact compute/memory footprints. These characteristics align naturally with the demands of energy-constrained or near-memory execution environments, where classical deep learning often proves too heavy [1, 5, 28, 16, 22].
Figure 1: Our proposed prototype reduction strategy. Conventional models store one dense hypervector per class ($\mathcal{O}(CD)$). DecoHD composes prototypes from a small shared set of channels, yielding $M=\prod_{i}L_{i}$ bound paths with $\sum_{i}L_{i}\leq M\leq C$, reducing memory while preserving full-$D$ representations. As illustrated in the bottom-left figure, varying the number of layers $N$ and the channel count $\sum_{i}L_{i}$ enables a tunable trade-off among memory footprint, inference cost, accuracy, and robustness, whereas prior feature-axis methods collapse to a single point at fixed dimensionality.
A conventional HDC classifier stores one prototype hypervector per class and predicts by measuring similarity between a query and these prototypes, typically via dot product [9]. While this layout is simple and hardware-friendly, its storage requirement grows as $\mathcal{O}(CD)$ for $C$ classes and dimensionality $D$, which quickly dominates memory in multi-class problems. Prior work has largely reduced this footprint by compressing along the feature axis: lowering $D$, imposing sparsity, or quantizing encodings [13, 24, 8, 31, 11]. Although effective in reducing memory and multiply–accumulate cost, such strategies erode the concentration-of-measure properties that stabilize similarity in high dimensions. At extreme budgets, they degrade both robustness to analog noise [18, 19] and classification accuracy.
This tension is especially acute in deployment scenarios where memory is the first-order constraint. Across emerging AI-on-edge applications—such as TinyML-based sensing and recognition [2, 17, 20], batteryless IoT platforms [5, 29], and federated/distributed AI in heterogeneous systems [6, 40, 3]—models must be aggressively compressed while retaining resilience to device-level imperfections. In these regimes, preserving $D$ maintains the concentration-of-measure properties central to HDC, making class-axis compression a more principled pathway than shrinking dimensionality.
Inspired by decomposition techniques successful in compact deep neural networks—such as low-rank adaptation, modular fine-tuning, and shared-codebook representations [34, 33, 10, 30, 15]—we propose a complementary direction tailored to HDC. Rather than compressing hypervector dimensionality, we reduce the number of stored dense hypervectors by parameterizing class prototypes as structured compositions of a small, shared set of channels, combined using HDC-native binding ($\otimes$) and weighted bundling ($\oplus$), as illustrated in Figure 1. Concretely, $N$ layers expose $\{L_{i}\}$ learnable channels; selecting one channel per layer yields $M=\prod_{i}L_{i}$ bound paths. A lightweight class-specific head $W\in\mathbb{R}^{C\times M}$ aggregates these paths to form logits. This design scales memory roughly with $\sum_{i}L_{i}$, while maintaining full-$D$ representations and their robustness benefits. This decomposition also introduces a tunable memory–compute trade-off: increasing the number of layers $N$ or channels reduces storage but raises test-time computation, since more bound-path compositions must be evaluated; deeper factorizations with small $L$ provide larger representational capacity at the cost of higher inference cost, a favorable exchange in regimes where memory is the primary constraint.
Naively decomposing dense prototypes post hoc is unstable in high dimensions, since binding is multiplicative and orderless, making the inverse factorization ill-posed [7, 25]. Instead, we introduce DecoHD, which learns directly in the decomposed space. Low-dimensional latents are expanded into channel hypervectors via fixed random projections, and training optimizes the bind–bundle pipeline end-to-end using cross-entropy. At inference, the method relies solely on $\otimes$, $\oplus$, and dot products, ensuring compatibility with existing HDC/VSA accelerators and memory-centric hardware paths.
In summary, DecoHD tackles the memory bottleneck of conventional HDC by decomposing class prototypes into a compact set of shared channels while preserving the ambient dimensionality $D$. This design retains the stability and robustness of high-dimensional representations under extreme budgets and aligns with the needs of edge-AI deployments across TinyML, federated learning, and memory-centric accelerators, where compact yet reliable models are critical [16]. Our evaluation shows that DecoHD achieves accuracy comparable to conventional HDC while delivering up to $277\times$ higher energy efficiency and $35\times$ faster inference than CPU baselines, $14\times$ and $3.7\times$ improvements over GPU, and $2.0\times$ and $2.4\times$ gains over a conventional HDC ASIC, all with only $0.38\times$ memory.
The main contributions of this work are:
  1. 1.
    Decomposed HDC classifier. We propose DecoHD, which replaces $C$ dense prototypes with a compact set of shared channels and a lightweight bundling head. This reduces memory from $\mathcal{O}(CD)$ to $\mathcal{O}(LD)$ with $L\!\ll\!C$, while preserving dimensionality $D$ and robustness to device-level noise.
  2. 2.
    End-to-end HDC-aware training. We develop a bind–bundle pipeline that jointly optimizes low-dimensional latents (expanded via fixed random projections) and class bundling weights. Unlike post hoc factorization, this preserves holographic properties of HDC and enables stable learning directly in decomposed space, reducing trainable parameters by up to 97%.
II Related Work
II-A HDC and Vector Symbolic Architectures
Hyperdimensional computing—also known as vector symbolic architectures—represents symbols and compositional structure with high-dimensional hypervectors manipulated by binding ($\otimes$), bundling ($\oplus$), and permutation ($\pi$[14]. Classification typically superposes encoded examples into per-class prototypes and scores by dot product. This operation set maps well to emerging in/near-memory platforms and low-power edge contexts: recent work demonstrates robust HDC pipelines under voltage scaling and binary encodings [19], specialized VSA/HDC designs for efficiency [23], and accelerator paths that preserve the native HDC interface [22]. At the system level, the push toward consumer edge-AI [16] and intermittently powered/batteryless devices [1, 5, 29] further motivates lightweight, memory-frugal models. Concurrently, the embedded-systems community highlights in/near-memory directions [28] and 3D integration trends [34, 33] that favor linear, data-parallel HDC kernels. Our design intentionally keeps HDC’s native bind–bundle–dot scoring to remain drop-in compatible with such accelerators.
II-B Model-Size Reduction for HDC-based Classifiers
Many efficiency efforts in HDC implicitly shrink along the feature axis—e.g., binary/low-precision encodings or simplified encoders—which reduce stored bits and arithmetic but also reduce effective dimensionality, weakening high-$D$ averaging that stabilizes similarity and increasing sensitivity to perturbations under tight power/voltage budgets [19, 39]. In contrast, DecoHD preserves the ambient $D$ and compresses orthogonally along the class axis by sharing a small pool of per-layer channels across classes and learning only the light bundling head. This reduces the # of dense hypervectors without altering the bind–bundle–score pipeline, and complements systems work on efficient HDC datapaths and accelerators [22, 23].
II-C Decomposition in DNNs and HDC for Model Compression
Decomposition is a standard route to shrink DNNs while preserving input/output dimensionality—e.g., trainable low-rank/tensor factorizations and related adapters [10, 30]. Our approach brings this spirit to HDC but avoids ill-posed post hoc factorization of dense class prototypes. Instead, DecoHD learns in a decomposed parameterization: a small, shared set of per-layer channels whose stacked bindings yield $M$ path features, followed by class-wise weighted bundling. This keeps $D$ unchanged, retains native HDC operations ($\otimes,\oplus$), and shifts storage from $C\times D$ prototypes to compact channels plus a light bundling head—well-suited to near-memory execution and resource-constrained edge deployment [28, 16, 1]. Finally, our end-to-end training of low-dimensional latents (expanded by fixed random projections) aligns with recent trends in efficient/edge training and personalization [15, 6, 40].
Figure 2: DecoHD overview. A fixed encoder maps inputs to hypervectors $h\!\in\!\mathbb{R}^{D}$. NN layers each provide $L_{i}$ learnable channels $\{A^{(i)}_{\ell}\}$ generated from low-dimensional latents via frozen projections. For an input, all $M=\prod_{i}L_{i}$ path hypervectors are formed by successive binding ($\otimes$); class-wise vectors are then produced by weighted bundling ($\oplus$) with $W\!\in\!\mathbb{R}^{C\times M}$ and scored against $h$ via dot products. Training updates only the latents and $W$.
III Methodology
Input : Data $\mathcal{D}=\{(x,y)\}$, fixed encoder $\phi:\mathbb{R}^{d_{\text{in}}}\!\to\!\mathbb{R}^{D}$, layers NN, channels $\{L_{i}\}_{i=1}^{N}$, latent dim $d$, fixed projectors $R^{(i)}\!\in\!\mathbb{R}^{d\times D}$.
Output : Latents $\{a^{(i)}_{\ell}\!\in\!\mathbb{R}^{d}\}$ and bundling weights $W\!\in\!\mathbb{R}^{C\times M}$ with $M=\prod_{i=1}^{N}L_{i}$.
Init:
 $a^{(i)}_{\ell}\!\sim\!\mathcal{N}(0,\sigma^{2})$,
  $W_{c,m}\!\leftarrow\!1/M$;
  freeze $\phi$ and all $R^{(i)}$.
Training (for epochs):
  1. 1.
    Sample mini-batch $\{(x_{b},y_{b})\}_{b=1}^{B}$; encode $h_{b}\!\leftarrow\!\phi(x_{b})$.
  2. 2.
    Compose channels: $A^{(i)}_{\ell}\!\leftarrow\!a^{(i)}_{\ell}R^{(i)}\in\mathbb{R}^{D}$ for all $i,\ell$.
  3. 3.
    For each $b$: form all path HVs by stacked binding $Z_{m}(h_{b})\leftarrow h_{b}\otimes\bigotimes_{i=1}^{N}A^{(i)}_{m_{i}}$ for m=1..Mm\!=\!1..M.
  4. 4.
    Bundle per class and score: $Y_{c}(h_{b})\!\leftarrow\!\bigoplus_{m=1}^{M}W_{c,m}Z_{m}(h_{b})$,
$s_{c}(x_{b})\!\leftarrow\!\langle Y_{c}(h_{b}),\,h_{b}\rangle$.
  • 5.
    Loss: $\mathcal{L}\!=\!\frac{1}{B}\sum_{b}\!-\log\frac{e^{s_{y_{b}}(x_{b})}}{\sum_{c}e^{s_{c}(x_{b})}}$;
  •   update $\{a^{(i)}_{\ell}\}$ and $W$ with AdamW.
[…]
    1. 1.
      $h\!\leftarrow\!\phi(x_{q})$;
    materialize $\{A^{(i)}_{\ell}\}$.
  • 2.
    Compute $Z_{m}(h)$ sequantially;
  •   $Y_{c}(h)\!\leftarrow\!\bigoplus_{m}W_{c,m}Z_{m}(h)$;
    $s_{c}\!\leftarrow\!\langle Y_{c}(h),h\rangle$.
  • 3.
    return $\hat{y}\!=\!\arg\max_{c}s_{c}$.
  • Algorithm 1 DecoHD: Training & Inference
    Conventional HDC stores one $D$-dimensional prototype per class, yielding $\mathcal{O}(CD)$ memory. Directly decomposing trained class hypervectors into a small set of shared factors is ill-posed in HD spaces: binding is multiplicative and orderless, and many distinct factorizations yield near-identical similarities. DecoHD sidesteps this by learning directly in a decomposed parameterization as illustrated in Figure 2. We introduce $N$ layers; layer $i$ exposes $L_{i}$ learnable channels—each channel is a factor hypervector that can bind with the input. Selecting one channel per layer defines a path; all $M=\prod_{i=1}^{N}L_{i}$ paths are formed by stacked binding and then bundled with class-specific weights to produce class vectors. This turns a small linear budget in $\sum_{i}L_{i}$ channels into a combinatorial set of $M$ bound paths while keeping the HD operations native. Detailed procedures for each stage is described in Algorithm 1.
    III-A Preliminaries
    Let $D$ denote the hypervector dimension and $C$ the number of classes. HDC represents items as $D$-dimensional real hypervectors and uses two primitives: binding ($\otimes$) as elementwise multiplication and bundling ($\oplus$) as (weighted) elementwise addition. An encoder $\phi:\mathbb{R}^{d_{\mathrm{in}}}\!\rightarrow\!\mathbb{R}^{D}$ maps inputs to hypervectors; we use a fixed random projection $\phi(x)=xW_{\mathrm{enc}}$ (Gaussian or ternary). Similarity is measured by the dot product, and we train with cross-entropy.
    We write $N$ for the number of layers, $L_{i}$ for the number of channels in layer $i$, and
    $M\;=\;\prod_{i=1}^{N}L_{i}$ (1)
    for the number of binding paths. The class-wise bundling weights are $W\!\in\!\mathbb{R}^{C\times M}$.
    III-B DecoHD Train
    Channels from low-dimensional latents. Layer $i$ contains $L_{i}$ low-dimensional latents $a^{(i)}_{\ell}\!\in\!\mathbb{R}^{d}$ with $d\!\ll\!D$. Each latent is expanded by a frozen random projector $R^{(i)}\!\in\!\mathbb{R}^{d\times D}$ to a channel (factor) hypervector:
    $A^{(i)}_{\ell}\;=\;a^{(i)}_{\ell}\,R^{(i)}\;\in\;\mathbb{R}^{D}.$ (2)
    Only the latents $\{a^{(i)}_{\ell}\}$ and the class bundling weights $W$ are learned; all $R^{(i)}$ and the input encoder $W_{\mathrm{enc}}$ remain fixed.
    Path composition by stacked binding. Given $h=\phi(x)$, picking one channel per layer yields a path $m=(m_{1},\dots,m_{N})\in[L_{1}]\!\times\!\cdots\!\times\![L_{N}]$. We construct the path hypervector by binding $h$ with the selected channels:
    $Z_{m}(h)\;=\;h\;\otimes\;\bigotimes_{i=1}^{N}A^{(i)}_{m_{i}}\;\in\;\mathbb{R}^{D}.$ (3)
    Broadcasted binding realizes all $M$ paths layer-by-layer.
    Class bundling and logits. Class $c$ aggregates path hypervectors with learned weights $W_{c,m}$:
    $Y_{c}(h)\;=\;\bigoplus_{m=1}^{M}W_{c,m}\,Z_{m}(h)\;\in\;\mathbb{R}^{D},$ (4)
    and we compute dot-product logits
    $s_{c}(x)\;=\;\langle Y_{c}(h),\,h\rangle.$ (5)
    Training. During training, we minimize cross-entropy over logits Equation 5:
    $\mathcal{L}_{\mathrm{CE}}(x,y)\;=\;-\log\frac{\exp\big(s_{y}(x)\big)}{\sum_{c=1}^{C}\exp\big(s_{c}(x)\big)}.$ (6)
    Gradients backpropagate through the $\otimes\!\rightarrow\!\oplus$ pipeline to the latents $\{a^{(i)}_{\ell}\}$ and $W$. Keeping $R^{(i)}$ and $W_{\mathrm{enc}}$ fixed (i) preserves holography in the HD space and (ii) makes optimization efficient because the learnable parameters reside in the low-dimensional latent space.
    III-C DecoHD Inference
    At test time we materialize the channels $\{A^{(i)}_{\ell}\}$ once from the trained latents and perform sequential test-time composition under extreme memory budgets. Rather than materializing all $M$ path features, we stream them one-by-one:
    $\underbrace{\text{stacked binding}}_{\otimes}$ $\;\Rightarrow\;\underbrace{\text{class bundling (accumulate)}}_{\oplus}$
    $\;\Rightarrow\;\underbrace{\text{dot-product scoring}}_{\langle\cdot,\cdot\rangle}.$
    Concretely, we iterate over $m=1,\dots,M$: (i) form $Z_{m}(h)$ via Equation 3 by sequentially binding the selected channels; (ii) immediately accumulate into class bundles $Y_{c}\!\leftarrow\!Y_{c}\oplus W_{c,m}Z_{m}(h)$ via Equation 4; after the sweep, (iii) compute logits $s_{c}(h)=\langle Y_{c}(h),h\rangle$ (Equation 5) and return $\arg\max_{c}s_{c}(x)$. All operations stay in $\mathbb{R}^{D}$ and use native HDC primitives. We can further shrink peak memory by conducting score-only streaming, skipping storing $Y_{c}$ and updating scores directly as $s_{c}\!\leftarrow\!s_{c}+W_{c,m}\,\langle Z_{m}(h),h\rangle$, discarding $Z_{m}(h)$ after each step.
    III-D Memory and Compute
    A conventional HDC table stores $C$ prototypes, i.e., $\mathcal{O}(CD)$ reals. DecoHD stores (i) $\sum_{i=1}^{N}L_{i}$ shared channels in $\mathbb{R}^{D}$ (or, during training, only low-dimensional latents plus fixed $R^{(i)}$), and (ii) a lightweight bundling head $W\!\in\!\mathbb{R}^{C\times M}$ with $M=\prod_{i}L_{i}$. Per sample, we use only native HDC primitives, but the arithmetic differs from a prototype table (which needs only $C$ dot products). In the streaming regime we iterate over paths $m=1,\dots,M$: (i) form $Z_{m}(h)$ by $N$ element-wise bindings, (ii) compute $t_{m}=\langle Z_{m}(h),h\rangle$, and (iii) update scores $s_{c}\leftarrow s_{c}+W_{c,m}\,t_{m}$. Streaming keeps peak memory to one working hypervector ($\mathcal{O}(D)$) plus $C$ scalars, trading extra per-sample work for a much smaller footprint than storing $C$ full prototypes. If memory permits pre-materializing class prototypes $Y_{c}$, inference reduces to the conventional $C$ dot products.
    IV Experiments
    TABLE I: Datasets used in evaluations. $C$ denotes the number of classes.
    Dataset # Features $\boldsymbol{C}$ # Train # Test Description
    ISOLET [4] 617 26 6,238 1,559 Voice recognition
    UCIHAR [27] 261 12 6,213 1,554 Activity recognition (mobile)
    PAMAP2 [26] 75 5 611,142 101,582 Activity recognition (IMU)
    PAGE [21] 10 5 4,925 548 Page layout blocks classification
    Figure 3: Accuracy versus a strong HDC baseline. Values in parentheses denote the target memory budget $m$ enforced relative to a conventional prototype table. We report test accuracy across numeric precisions and hypervector dimensionalities $D$, and compare to OnlineHD (no parameter reduction).
    Figure 4: Accuracy compared to state-of-the-art feature-axis compression. Values in parentheses denote the target memory budget $m$. To isolate the impact of class-axis decomposition, we compare against SparseHD, a representative feature-axis reduction method, under matched budgets.
    Figure 5: Robustness to random bit-flip noise. Values in parentheses denote the target memory budget $m$. We inject independent random bit flips into 32-bit floating-point representations and evaluate accuracy as the flip probability increases.
    Figure 6: Sensitivity to latent dimensionality. We evaluate a two-layer $\lfloor\sqrt{C}\rfloor\times\lfloor\sqrt{C}\rfloor$ configuration with $d\in\{256,1024,4096\}$ at $D\in\{1{,}000,10{,}000\}$ and report test accuracy.
    Figure 7: Sensitivity to the number of layers. We vary the number of layers ($N\in\{1,2,3\}$) under multiple target memory budgets, with up to five channels per layer, and report test accuracy for $d\in\{256,4096\}$. The dotted curve is a logarithmic fit for visualization.
    IV-A Experimental Setup
    Implementation and precision. All models are implemented in PyTorch with a precision-aware training loop. Unless otherwise stated, training uses 32-bit floating point with AdamW. We additionally evaluate reduced-precision execution via native fp16/bf16 (AMP on GPUs and bf16 on CPU) and emulated fp8/fp4 by rounding mantissa/exponent bits. When available, we also include CUDA fp8_e4m3fn/fp8_e5m2. No integer quantization or post-training calibration is applied.
    Datasets and preprocessing. We follow the standard train/test splits for ISOLET, UCIHAR, PAGE, and PAMAP2 (see Table I for details). Features are standardized to zero mean and unit variance using statistics from the training split only, and all results are reported on the held-out test split.
    Encoder. Inputs are mapped to $D$-dimensional real hypervectors using a fixed random encoder $\phi(x){=}xW_{\text{enc}}$ (Gaussian by default; ternary ablations included). Unless noted, we use D=10, [^0] D{=}10{,}000 and d=4,096d{=}4{,}096.
    Our model (DecoHD). We train directly in a decomposed HDC parameterization with $N\in\{1,2,3\}$ layers and per-layer channel counts $L_{i}\in\{1,\dots,5\}$ (grid over all nonempty tuples). Each channel is generated from a low-dimensional latent (default d=4096d{=}4096) through a frozen random projection into $\mathbb{R}^{D}$. Given an encoded input $h$, stacked binding ($\otimes$) across layers yields $M{=}\prod_{i}L_{i}$ path hypervectors that are aggregated by a lightweight class bundling head $W\in\mathbb{R}^{C\times M}$ via $\oplus$. Logits are dot products with $h$; we train with cross-entropy. Optimization uses AdamW (learning rate $1\!\times\!10^{-3}$, weight decay $5\!\times\!10^{-5}$), batch size $1024$ with microbatches of $128$, for $1000$ epochs. $W$ is initialized uniformly to $1/M$; the input and per-layer projections remain fixed. Inference follows the same bind $\rightarrow$ bundle $\rightarrow$ dot pipeline.
    Baselines. We compare against (i) a conventional HDC prototype table constructed by class-wise summation and (ii) OnlineHD with iterative refinement for $200$ epochs (learning rate $0.1$). For feature-axis reduction, we include SparseHD under matched memory budgets. All baselines share the same encoder $\phi$ and preprocessing and are evaluated under identical precision settings.
    Target memory budgets. Results are reported under budgets $\leq m\in(0,1]$ and denoted in parentheses as $(\leq m)$, where $m$ is the model size relative to a conventional HDC table with $C\times D$ parameters (at $b$ bits each). DecoHD replaces this table with a lightweight bundling head $W\!\in\!\mathbb{R}^{C\times M}$ and a bank of channel latents totaling $L_{\mathrm{tot}}=\sum_{i}L_{i}$ of size $d$, yielding the approximate normalized footprint $m\;\approx\;\frac{CM+L_{\mathrm{tot}}D}{CD},$ with $M=\prod_{i}L_{i}$. Selecting $(N,\{L_{i}\},D)$ to satisfy $m$ enforces the target budget.
    IV-B Performance of DecoHD
    Figure 3 compares DecoHD to the strong non-reduced HDC baseline, OnlineHD, across numeric precisions and hypervector dimensionalities $D$ under target memory budgets $m$ (values shown in parentheses). Despite operating with substantially fewer trainable parameters, DecoHD tracks OnlineHD closely across datasets and settings. At $D{=}10\mathrm{K}$, the average accuracy gaps are small—approximately $0.15\%$ at $m{\leq}0.7$ and $0.1\%$ at $m{\leq}0.5$—with a worst case of about $5.7\%$ even at the tight $m{\leq}0.5$ budget. When both precision and dimensionality are reduced, the gap grows modestly as expected (e.g., roughly $0.7\%$ at 16-bit and $D{=}1\mathrm{K}$ for $m{\leq}0.5$), yet remains minor in absolute terms. Two trends recur across settings. First, lower numeric precision slightly increases the gap to OnlineHD but does not alter the qualitative ranking, indicating that the decomposition does not introduce precision-specific failure modes. Second, lowering $D$ reduces the usual concentration-of-measure benefits of HDC; nevertheless, DecoHD remains competitive, suggesting that composing class prototypes from a shared channel bank preserves much of the structure that OnlineHD leverages at full size. Overall, across budgets, precisions, and dimensionalities, DecoHD delivers accuracy within a tight margin of OnlineHD while meeting strict memory targets, demonstrating that training directly in the decomposed parameterization is an effective route to compress HDC models with negligible loss.
    IV-C Comparison to feature-axis reduction
    Figure 4 contrasts DecoHD with SparseHD, a state-of-the-art feature-axis reduction method that compresses by shrinking $D$. Under matched budgets, DecoHD outperforms SparseHD in all cases for $m{\leq}0.7$ and in six of eight cases for $m{\leq}0.5$. The difference reflects where compression is applied. Feature-axis reduction weakens high-dimensional orthogonality and the concentration properties central to HDC’s separability and robustness; as $D$ is reduced, prototype quality and similarity estimates degrade, and the classifier becomes more sensitive to noise and incidental correlations. DecoHD, by contrast, preserves the full ambient dimensionality and instead compresses along the class axis via decomposition, sharing a compact set of expressive channels across classes using binding and bundling. The class head $W$ then allocates these shared channels adaptively. This strategy is more resilient at tight budgets because it amortizes parameters over classes while retaining the geometric advantages of a large $D$.
    IV-D Robustness of DecoHD
    Figure 5 evaluates tolerance to random bit-flip noise injected into 32-bit floating-point representations. At a fixed $D{=}10\mathrm{K}$, DecoHD maintains higher accuracy than OnlineHD as the flip probability increases, indicating that the decomposition does not merely match clean accuracy but also confers stability under perturbation. The effect can be understood through two complementary mechanisms that operate within the same encoder and hyperspace. First, bundling averages over multiple bound channels, so perturbations that affect individual channels tend to be attenuated when aggregated, reducing the variance of the effective class representation encountered at inference. Second, binding behaves like a quasi-orthogonalizing transformation; independent bit flips in one channel yield perturbations that are largely decorrelated from those in other channels, limiting coherent error accumulation. Since both methods share the same input encoding and dimensionality, the robustness margin is attributable to DecoHD’s representation strategy rather than differences in $D$ or preprocessing, which is particularly relevant for near-memory or in-sensor deployments where soft errors and read-disturb effects are non-negligible.
    IV-E Impact of latent dimensionality
    Figure 6 studies latent sizes $d\in\{256,1024,4096\}$ at $D\in\{1{,}000,10{,}000\}$ using a two-layer $\lfloor\sqrt{C}\rfloor\times\lfloor\sqrt{C}\rfloor$ configuration. Accuracy generally saturates by d=4096d{=}4096 across datasets, indicating that channel latents need not match the ambient dimension to achieve strong performance. Tasks with fewer classes, such as PAGE and PAMAP2, reach their accuracy plateau as early as d=256d{=}256, reflecting lower intrinsic class complexity and correspondingly lighter demands on channel expressivity. Because the latent bank scales with $L_{\mathrm{tot}}d$, choosing $d\!\ll\!D$ yields substantial parameter savings: in our settings, up to $97.44\%$ fewer trainable parameters compared to directly learning a full $C{\times}D$ prototype table, with minimal or no loss once the dataset-specific plateau is reached. A practical guideline is to select the smallest $d$ on the observed saturation plateau for the target dataset; combined with a modest $M$, this typically meets the memory target $m$ while preserving accuracy.
    IV-F Impact of the number of layers
    Figure 7 varies the number of layers $N\in\{1,2,3\}$ (with up to five channels per layer) across memory budgets for $d\in\{256,4096\}$. The results reveal a clear interaction between latent capacity and architectural depth. When $d$ is large ($4096$), fewer layers often achieve the best accuracy at a fixed budget, because channels are already expressive and additional binding primarily increases MM—and thus the size of WW—without commensurate gains. When $d$ is small ($256$), increasing $N$ is beneficial: deeper binding expands the combinatorial basis of path hypervectors and improves class separability at the same overall budget, effectively trading a modest increase in head size for a disproportionately large increase in representational diversity. From a budget-allocation standpoint, recall that $M=\prod_{i}L_{i}$ grows multiplicatively with $N$; increasing layers shifts budget from the latent bank ($L_{\mathrm{tot}}d$) toward the class head ($CM$). At small $d$, the marginal utility of enlarging $M$ is high, whereas at large $d$ it is lower because channels already encode rich structure. In practice, given a target $m$, it is effective to first pick the smallest $d$ that lies on the latent-dimension plateau (as identified in Figure 6); if this $d$ is small, favor $N\geq 2$ with moderate per-layer channel counts to grow $M$, whereas if $d$ is large, prefer $N\in\{1,2\}$ with slightly larger per-layer channel counts and allocate remaining budget to the class head. This recipe consistently stays within budget while preserving or improving accuracy.
    IV-G Hardware Efficiency and Trade-offs
    TABLE II: System-level efficiency of 1-layer DecoHD (ASIC) vs. a conventional HDC baseline executed on CPU/GPU/ASIC (ISOLET; D=10, [^0] D{=}10{,}000, C=26C{=}26, $L_{i}{=}10^{1/N}$).
    Platform Energy eff. ($\times$) $\uparrow$ Speedup ($\times$) $\uparrow$ Memory usage ($\times$) $\downarrow$
    CPU (Ryzen 9 9950X) 277.40 34.87 0.38
    GPU (RTX 4090) 13.53 3.66 0.38
    ASIC 2.02 2.42 0.38
    TABLE III: Depth trade-offs of DecoHD (ASIC) vs. a conventional HDC (ASIC) on ISOLET (D=10, [^0] D{=}10{,}000, C=26C{=}26, $L_{i}{=}10^{1/N}$).
    # Layers Speedup ($\times$) $\uparrow$ Memory usage ($\times$) $\downarrow$ Accuracy drop (%) $\downarrow$
    1-layer 2.42 0.38 1.01
    2-layer 0.94 0.12 9.81
    3-layer 0.89 0.08 28.7
    Table II shows that mapping DecoHD to an ASIC yields up to $277\times$ higher energy efficiency and strong speedups over CPU/GPU baselines, while remaining both faster and more energy-efficient than a conventional HDC ASIC, at only $0.38\times$ memory. As Table III indicates, depth induces a clear trade-off: a shallow (1-layer) design maximizes throughput with minimal accuracy loss; increasing depth further reduces memory but increases test time per-sample compute (reducing speed) and increases per-channel interference (reducing accuracy). Therefore, our findings suggest selecting the shallowest factorization that meets the memory budget to preserve throughput and accuracy; deeper configurations are justified only when memory is the primary system bottleneck.
    V Conclusions
    We presented DecoHD, a gradient-based class-axis decomposition for HDC that enables memory-lean inference under extreme resource constraints. Across benchmarks, it closely matches a strong non-reduced HDC baseline (average accuracy gap $\approx$0.1–0.15%), improves robustness to random bit-flip noise, and reduces trainable parameters by up to $\sim$97% at saturation. In hardware, an ASIC realization achieves up to $277\times$ higher energy efficiency at $0.38\times$ memory, with consistent speedups over CPU/GPU and a conventional HDC ASIC. By preserving full hypervector dimensionality while aggressively reducing stored prototypes, DecoHD is well suited for TinyML and edge deployments in memory-first and near-memory systems. Depth reduces memory but increases test-time computation; thus, the shallowest factorization that meets the target memory budget is preferred to preserve throughput and accuracy.
    Acknowledgements
    This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2431561, #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360, the U.S. Army Combat Capabilities Development Command Army Research Laboratory under Support Agreement No. USMA 21050, and DARPA under Support Agreement No. USMA 23004. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.